OpenStudy (fairy733189):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t^2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is how many feet per second?

OpenStudy (fairy733189):

@math&ing001

Mehek (mehek14):

It's impossible to tag him :P

Mehek (mehek14):

Plug 4 for t and solve

OpenStudy (fairy733189):

I plugged 4 in for t and then 6 in for t because I know that's what you have to do but after that how do I solve?

Mehek (mehek14):

Can you tell me what you get when t = 4 and t = 6?

OpenStudy (fairy733189):

I got: f(4) = -16(4)^2 + 32(4) + 90 f(4) = -256 + 128 + 90 f(4) = 14 f(6) = -16(6)^2 + 32(6) + 90 f(6) = -512 + 192 + 90 f(6) = -230

OpenStudy (fairy733189):

@Mehek14

Mehek (mehek14):

I got 4, -38 :\ 6, -294

OpenStudy (fairy733189):

hmm how did you do the problems then?

Mehek (mehek14):

f(t) = -16*4^2 + 32 * 4 + 90 -16 * 16 + 128 + 90 -256 + 128 + 90 -38

Mehek (mehek14):

f(t) = -16*6^2 + 32*6 + 90 -576 + 192 + 90 -294

OpenStudy (fairy733189):

I see what I did wrong

OpenStudy (fairy733189):

So after that what am I supposed to do?

Mehek (mehek14):

Slope formula \(\bf{\dfrac{y_2-y_1}{x_2-x_1}}\)

Mehek (mehek14):

So plug in the values \(\bf{\dfrac{-294 - (-38)}{6-4}}\)

OpenStudy (fairy733189):

I got: -256/2 -128

Mehek (mehek14):

So -128 is the answer

OpenStudy (fairy733189):

but is it actually the answer because the ball can go -128 ft.

Mehek (mehek14):

Yes because it is dropped from a building

OpenStudy (fairy733189):

oh ok I understand. Thanks so much!!! Can you help me with another problem?

Mehek (mehek14):

Sure

OpenStudy (fairy733189):

do you want me to close this one first or can I just ask it here?

Mehek (mehek14):

your choice

OpenStudy (fairy733189):

I closed it and I tagged you in the other problem